# A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process

### v1

@@ 3.2 Injectivity of the map, $\mathcal{P}_{p}$
Although $\mathcal{P}_{p}$ in (13) defines a map from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$, it is important to guarantee that such a map is injective (i.e., a subspace in $\mathcal{G}(r,n)$ can be uniquely determined by a vector in $\Omega_{p}$). Proposition 3.2 below gives the condition for which $\mathcal{P}_{p}$ is injective.
Proposition 3.2.
Given a basepoint $p\in\mathcal{G}(r,n)$, there exists an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$
(15) $\Omega_{p}:=\left\{\bm{y}\in\mathbb{R}^{nr-r};||\bm{y}||<\frac{\pi}{2}\right\}$
@@
Proof. Consider an open set $D_{p}\subset\mathcal{T}_{p}\mathcal{G}(r,n)$
(16) $D_{p}:=\left\{v\in\mathcal{T}_{p}\mathcal{G}(r,n);||v||<\pi/2\right\}$
where $\pi/2$ is the injective radius on the Grassmann Manifold[12]. Hence, we seek an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$ such that $\mathcal{A}_{p}$ maps a vector $\bm{y}\in\Omega_{p}$ to the open disk $D_{p}$ above. Because $\bm{Z}=\text{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$, we have
(17) $\begin{split}\bm{Z}^{T}\bm{Z}&=\begin{bmatrix}\bm{y}^{T}\bm{F}_{1}\\ \bm{y}^{T}\bm{F}_{2}\\ \vdots\\ \bm{y}^{T}\bm{F}_{r}\end{bmatrix}\begin{bmatrix}\bm{F}_{1}^{T}\bm{y}&\bm{F}_{2}^{T}\bm{y}&\cdots&\bm{F}_{r}^{T}\bm{y}\end{bmatrix}\\ &=\begin{bmatrix}\sum_{i=1}^{n-1}y_{(i-1)\times r+1}^{2}&&&\\ &\sum_{i=1}^{n-1}y_{(i-1)\times r+2}^{2}&&\\ &&\ddots&\\ &&&\sum_{i=1}^{n-1}y_{(i-1)\times r+r}^{2}\end{bmatrix}\end{split}$
@@
(18) $||v||=\sqrt{\text{trace}(\bm{Z}^{T}\bm{Z})}=\left(\sum_{i=1}^{nr-r}y_{i}^{2}\right)^{\frac{1}{2}}=||\bm{y}||$
which needs to be smaller than $\pi/2$. Hence, we have an open set $\Omega_{p}$, defined in (15), such that $\mathcal{P}_{p}$ is injective.
3.3 Prediction based on the pGP regression
Proposition 3.1 suggests an effective approach to establish the mapping between the parameter and POD basis. In a nutshell, as illustrated in Figure 3, a GP can be established to describe the mapping between the parameters space $\mathbb{R}^{d}$ and $\mathbb{R}^{nr-r}$, and then, the GP is projected to the Grassmann Manifold. As a result, for any new parameter $\bm{\theta}^{*}$, a vector $\bm{y}^{*}$ can be found and mapped back to the Grassmann Manifold to determine the optimal subspace onto which the full-order system is projected (equivalently, the POD basis matrix $\bm{\Phi}^{*}$).
@@ 3.3 Prediction based on the pGP regression
Let the vector $\bm{y}$ follows the multivariate Gaussian distribution, i.e., $\bm{y}\sim\mathcal{N}(\bm{0}_{nr-r,1},\bm{K})$ where $\bm{K}$ is a $(nr-r)\times(nr-r)$ symmetric and positive semi-definite matrix. Then, $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ follows a matrix-variate Gaussian, $\bm{Z}\sim\mathcal{N}(\bm{0}_{n\times r},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$ in the sense that $\text{vec}(\bm{Z})\sim\mathcal{N}(\bm{0}_{nr\times 1},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$. For a given point $\mu\in\mathcal{G}(r,n)$ and a symmetric positive semi-definite matrix $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, we say that a random point $X\in\mathcal{G}(r,n)$ follows a Projected Gaussian Distribution (PGD) on $\mathcal{G}(r,n)$
(19) $X\sim\mathcal{P}_{\mu}(\mathcal{N}(\bm{0}_{nr-r,1},\bm{K}))$
@@
(20) $\mathcal{P}_{\mu}:=\text{span}\circ\mathcal{P}_{2}\circ\mathcal{P}_{1}$
@@
Proposition 3.3.
If $X_{i}\sim\mathcal{N}_{\mathcal{G}_{i}}(\mu_{i},\bm{K}_{i})$ for $i=1,2$, then, $(X_{1},X_{2})$ are jointly PGD on $\mathcal{G}_{1}\times\mathcal{G}_{2}$
(21) $(X_{1},X_{2})\sim\mathcal{N}_{\mathcal{G}_{1}\times\mathcal{G}_{2}}\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\tilde{\bm{K}}\right),\quad\quad\tilde{\bm{K}}=\begin{pmatrix}\bm{K}_{1},\bm{K}_{12}\\ \bm{K}_{12}^{T},\bm{K}_{2}\end{pmatrix}$
@@
Proposition 3.3 is obtained following a similar technique in [21]. Let $\mathbb{A}_{1}=\mathcal{P}_{\mu_{1}}^{-1}(p_{1})=\{\bm{y}\in\mathbb{R}^{nr-r}:\mathcal{P}_{\mu_{1}}(\bm{Y})=p_{1}\}$ be the preimage of $p_{1}$ in $\mathbb{R}^{nr-r}$ associated with $\bm{\Phi}_{\mu_{1}}$ given a point $\mu_{1}$, and let $\mathbb{A}_{2}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$ similarly be the preimage of $p_{2}$, then,
(23) $\begin{split}\text{Pr}(X_{1}|X_{2}=p_{2})&=\frac{\text{Pr}(\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2})}{\text{Pr}(\bm{y}_{2}\in\mathbb{A}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\frac{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}{\text{Pr}(\mathbb{A}_{2})}\frac{\mathcal{MN}((\bm{y}_{1}^{T},\bm{y}_{2}^{T})^{T};\bm{0},\tilde{\bm{K}})}{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\lambda_{\bm{y}_{2}}\mathcal{MN}(\bm{y}_{1};\bm{u}_{\bm{y}_{2}},\bm{K}_{\bm{y}_{2}})\end{split}$
@@
Once a PGD is obtained as described above, a Projected Gaussian Process (pGP) on Grassmann Manifold can be defined. We start with a multivariate GP defined in the Euclidean space. A collection $\bm{y}$ of random vectors is a Multivariate GP (MGP) in $\mathbb{R}^{d}$, denoted by $\bm{y}\sim\mathcal{MGP}(f,\omega,\bm{K})$, with a matrix-valued mean function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{nr-r}$, kernel $\omega:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, if any finite collection of $\bm{y}_{1}$, $\bm{y}_{2}$, $\cdots$, $\bm{y}_{k}$ have a joint multivariate Gaussian distribution such that
(25) $[\bm{y}_{1}^{T},\bm{y}_{2}^{T},\cdot,\bm{y}_{k}^{T}]^{T}\sim\mathcal{N}([f^{T}(\bm{\theta}_{1}),f^{T}(\bm{\theta}_{2}),\cdot,f^{T}(\bm{\theta}_{k})^{T}]^{T},\bm{\Omega}\otimes\bm{K})$
where $\bm{\Omega}\in\mathbb{R}^{k\times k}$ with $\bm{\Omega}_{i,j}=\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$. Then, a pGP on $\mathcal{G}(r,n)$ can be defined as follows:
Projected Gaussian Process. A Projected Gaussian Process (pGP) on $\mathcal{G}(r,n)$ is defined by
(26) $g\sim\mathcal{PGP}(m,\omega,\bm{K})$
@@
(27) $\begin{pmatrix}p_{*}\\ \bm{p}^{\text{train}}\end{pmatrix}\sim\mathcal{N}_{\mathcal{G}_{*}\times\mathcal{G}_{1}\times\cdots\times\mathcal{G}_{k}}\left(\begin{pmatrix}m_{*}\\ \bm{m}^{\text{train}}\end{pmatrix},\begin{pmatrix}\bm{K}_{**},\bm{K}_{*}\\ \bm{K}_{*}^{T},\bm{K}^{\text{train}}\end{pmatrix}\right)$
@@
(28) $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$
@@
(29) $\bm{u}_{*}=\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}((\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T},(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T},\cdots,(\mathcal{P}_{m(\bm{\theta}_{d})}^{-1}(p_{k}))^{T})^{T}$
@@
(30) $\tilde{\bm{K}}_{*}=\bm{K}_{**}-\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}\bm{K}_{*}^{T}.$
Hence, a point estimate of the subspace $p_{*}\in\mathcal{G}(r,n)$ can be obtained by $\hat{p}_{*}=\mathcal{P}_{m_{*}}(\bm{u}_{*})$. Because $\bm{u}_{*}$ is the mean of the multivariate Gaussian distribution $\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$ defined in $\mathcal{R}^{nr-r}$, $\hat{p}_{*}$ is a maximum a posteriori probability (MAP) estimate. Also recall that $||\bm{u}_{*}||$ needs to be smaller than $\pi/2$ for $\mathcal{P}_{m_{*}}$ to be injective. Hence, if $||\bm{u}_{*}||>\pi/2$, we shrink this vector $\bm{u}_{*}\leftarrow\frac{\pi}{2}\frac{\bm{u}_{*}}{||\bm{u}_{*}||}$, which is the projection of the original vector $\bm{u}_{*}$ onto the surface of a $(nr-r)$-dimensional sphere. Similarly, the estimate $\hat{\bm{\Phi}}_{*}$ is obtained as $\mathcal{P}_{2}\circ\mathcal{P}_{1}(\bm{u}_{*})$ such that $\hat{p}_{*}$ is spanned by $\hat{\bm{\Phi}}_{*}$.
@@
Figure 4 illustrate the process of predicting the optimal subspace using pGP. In this illustration, suppose that we obtain the snapshot data from three parameter settings $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$, as well as the optimal subspaces $p_{1}$, $p_{2}$ and $p_{3}$ that can be found using POD. The prediction of optimal subspace involves the following steps:
@@
(iii) The covariance matrix $\tilde{\bm{K}}_{*}$ is computed on $\mathbb{R}^{nr-r}$, which is exactly the same as the convectional GP. Steps (ii) and (iii) together yield a multivariate Gaussian distribution of $\bm{y}_{*}$) on $\mathbb{R}^{nr-r}$, i.e., $\bm{y}_{*}\sim\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$.
(iv) Finally, the Gaussian distribution of $\bm{y}_{*}$ on $\mathbb{R}^{nr-r}$ (or, the predicted column vector $\bm{u}_{*}$ if only a point prediction is needed), is projected to the Grassmann Manifold through the projector $\mathcal{P}_{m(\bm{\theta}_{*})}$.
@@ 3.4 Discussions on model specification and parameter estimation
A close examination of (28) reveals an important difference between how prediction is done by conventional GP and the proposed pGP. Using a conventional GP, the point predictor is a linear combination of the observations and the weights are determined by the covariance function (i.e., the point prediction does not depend the mean of the GP; a property well-known in the literature). For the proposed pGP, as indicated by (28), the point prediction does depend on the function $m(\bm{\theta})$ that describes the relationship between the parameter and basepoint on the Grassmann Manifold. In other words, the choice of $m(\bm{\theta})$ determines the mapping $\mathcal{P}_{m}$. Two approaches can be used to determine the function $m(\bm{\theta})$. The first approach is to fix the basepoint for all parameters, i.e., $m(\bm{\theta})=\mu$. which can be obtained using a global POD basis. The second approach is to use another regression model to determine $m(\bm{\theta})$, such as the tree-based method described in [18].
@@
In addition, because $n$ is usually large, a practical way to parameterize the $(nr-r)\times(nr-r)$ covariance matrix $\bm{K}$ is to let $\bm{K}=\sigma_{\bm{K}}\bm{I}$ and estimate $\sigma_{\bm{K}}$ from data. As for the kernel function $\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$, existing Gaussian, Exponential or other kernel functions can be used, and the parameter $\bm{\xi}$ can be estimated from data. Hence, once $m(\bm{\theta})$ is determined, the cross-validation-based approach can be used to estimate $(\sigma_{\bm{K}},\bm{\xi})$ by minimizing $\sum_{i=1}^{k}\delta_{i}$, where the cross validation error is given by $\delta_{i}=||\bm{D}(\bm{\theta}_{i})-\hat{\bm{\Phi}}_{-i}\hat{\bm{\Phi}}_{-i}^{T}\bm{D}(\bm{\theta}_{i})||_{F}^{2}$, and $\hat{\bm{\Phi}}_{-i}$ is the estimated $\bm{\Phi}_{i}$ excluding data $\bm{D}(\bm{\theta}_{i})$.
@@
Also note that, for systems with high-dimensional parameter space (i.e., relatively large $d$), it is often the case that not all parameters are equally important in terms of how they affect the POD bases. In other words, the ROM can be sensitive to some parameters while independent of the others. Hence, when constructing the statistical learning model that establishes the mapping from parameter space to Grassmann Manifold, it is necessary to identify those relevant parameters. For the pGP model, one can achieve this goal using the Automatic Relevance Determination (ARD)[28]. To illustrate the idea, consider the squared Exponential kernel given by
(31) $\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}$
@@
where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2},\cdots,\xi_{d+2}^{-2}\}$. Here, the parameters $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ play a role of characteristic length-scales for parameter $\theta_{1},\theta_{2},\cdots,\theta_{d}$. If the estimated length-scales are very large, the corresponding parameters have little effects on the prediction and are practically removed from the model.
@@
Finally, it is worth noting that the proposed pGP naturally involves the use of the global POD basis as its special case. If $\xi_{1}=0$ and the estimated $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ are all close to zero, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ becomes practically zero. In this case, (29) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$. If the global POD basis is chosen as the reference point (which spans a subspace in the Grassmann Manifold where the tangent space is attached to) and the basepoint function $\bm{m}$ is invariant with $\bm{\theta}$, then, any subspace, $p_{i}$, corresponding to the ROM at parameter $\bm{\theta}_{i}$ on the Grassmann Manifold can be seen as i.i.d. samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$. Hence, the point prediction, $\hat{p}_{*}$, is just the subspace spanned by the global POD basis.

### v2

@@ 3.2 Injectivity of the map, $\mathcal{P}_{p}$
Although $\mathcal{P}_{p}$ in (13) defines a map from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$, it is important to guarantee that such a map is injective. Proposition 3.2 below gives the condition for which $\mathcal{P}_{p}$ is injective.
Proposition 3.2. Given a basepoint $p\in\mathcal{G}(r,n)$, there exists an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$
[…]
@@
To prove Proposition 3.2, consider an open set $D_{p}\subset\mathcal{T}_{p}\mathcal{G}(r,n)$
[…]
where $\pi/2$ is the injective radius on the Grassmann manifold (Friderikos et al.,, 2020). Hence, we seek an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$ such that $\mathcal{A}_{p}$ maps a vector $\bm{y}\in\Omega_{p}$ to the open disk $D_{p}$ above. Because $\bm{Z}=\text{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$, we have
$\begin{split}\bm{Z}^{T}\bm{Z}&=\begin{bmatrix}\bm{y}^{T}\bm{F}_{1}\\ \bm{y}^{T}\bm{F}_{2}\\ \vdots\\ \bm{y}^{T}\bm{F}_{r}\end{bmatrix}\begin{bmatrix}\bm{F}_{1}^{T}\bm{y}&\bm{F}_{2}^{T}\bm{y}&\cdots&\bm{F}_{r}^{T}\bm{y}\end{bmatrix}\\ &=\begin{bmatrix}\sum_{i=1}^{n-1}y_{(i-1)\times r+1}^{2}&&&\\ &\sum_{i=1}^{n-1}y_{(i-1)\times r+2}^{2}&&\\ &&\ddots&\\ &&&\sum_{i=1}^{n-1}y_{(i-1)\times r+r}^{2}\end{bmatrix}\end{split}$ (17)
@@
[…]
which needs to be smaller than $\pi/2$ according to (16). Hence, we have an open set $\Omega_{p}$, defined in (15), such that $\mathcal{P}_{p}$ is injective.
3.3 Prediction using the pGP regression
Propositions 3.1 and 3.2 suggest an effective approach to establish the mapping between the parameter and POD basis. In a nutshell, as illustrated in Figure 3, a GP can be established to describe the mapping between the parameters space $\mathbb{R}^{d}$ and $\mathbb{R}^{nr-r}$, and then, the GP is projected to the Grassmann manifold. As a result, for any new parameter $\bm{\theta}^{*}$, a vector $\bm{y}^{*}$ can be found and mapped back to the Grassmann manifold to determine the optimal subspace onto which the full-order system is projected (equivalently, the POD basis matrix $\bm{\Phi}^{*}$).
Let the vector $\bm{y}$ follow a multivariate Gaussian distribution, i.e., $\bm{y}\sim\mathcal{N}(\bm{0}_{nr-r,1},\bm{K})$ where $\bm{K}$ is a $(nr-r)\times(nr-r)$ symmetric and positive semi-definite matrix. Then, $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ follows a matrix-variate Gaussian, $\bm{Z}\sim\mathcal{N}(\bm{0}_{n\times r},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$ in the sense that $\text{vec}(\bm{Z})\sim\mathcal{N}(\bm{0}_{nr\times 1},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$. For a given point $\mu\in\mathcal{G}(r,n)$ and a symmetric positive semi-definite matrix $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, we say that a random point $X\in\mathcal{G}(r,n)$ follows a projected Gaussian Distribution (pGD) on $\mathcal{G}(r,n)$
$X\sim\mathcal{P}_{\mu}(\mathcal{N}(\bm{0}_{nr-r,1},\bm{K}))$ (19)
@@ 3.3 Prediction using the pGP regression
$\mathcal{P}_{\mu}:=\text{span}\circ\mathcal{P}_{2}\circ\mathcal{P}_{1}$ (20)
@@
Proposition 3.3. If $X_{i}\sim\mathcal{N}_{\mathcal{G}_{i}}(\mu_{i},\bm{K}_{i})$ for $i=1,2$, then, $(X_{1},X_{2})$ are jointly pGD on $\mathcal{G}_{1}\times\mathcal{G}_{2}$
$(X_{1},X_{2})\sim\mathcal{N}_{\mathcal{G}_{1}\times\mathcal{G}_{2}}\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\tilde{\bm{K}}\right),\quad\quad\tilde{\bm{K}}=\begin{pmatrix}\bm{K}_{1},\bm{K}_{12}\\ \bm{K}_{12}^{T},\bm{K}_{2}\end{pmatrix}$ (21)
@@
Proposition 3.3 is obtained following a similar technique in Mallasto and Feragen, (2018). Let $\mathbb{A}_{1}=\mathcal{P}_{\mu_{1}}^{-1}(p_{1})=\{\bm{y}\in\mathbb{R}^{nr-r}:\mathcal{P}_{\mu_{1}}(\bm{Y})=p_{1}\}$ be the preimage of $p_{1}$ in $\mathbb{R}^{nr-r}$ associated with $\bm{\Phi}_{\mu_{1}}$ given a point $\mu_{1}$, and let $\mathbb{A}_{2}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$ similarly be the preimage of $p_{2}$, then,
$\begin{split}\text{Pr}(X_{1}|X_{2}=p_{2})&=\frac{\text{Pr}(\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2})}{\text{Pr}(\bm{y}_{2}\in\mathbb{A}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\frac{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}{\text{Pr}(\mathbb{A}_{2})}\frac{\mathcal{MN}((\bm{y}_{1}^{T},\bm{y}_{2}^{T})^{T};\bm{0},\tilde{\bm{K}})}{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\lambda_{\bm{y}_{2}}\mathcal{MN}(\bm{y}_{1};\bm{u}_{\bm{y}_{2}},\bm{K}_{\bm{y}_{2}})\end{split}$ (23)
@@
Once a pGD is obtained as described above, a pGP on Grassmann manifold can be defined. We start with a multivariate GP defined in the Euclidean space. A collection $\bm{y}$ of random vectors is a Multivariate GP (MGP) in $\mathbb{R}^{d}$, denoted by $\bm{y}\sim\mathcal{MGP}(f,\omega,\bm{K})$, with a matrix-valued mean function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{nr-r}$, kernel $\omega:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, if any finite collection of $\bm{y}_{1}$, $\bm{y}_{2}$, $\cdots$, $\bm{y}_{k}$ have a joint multivariate Gaussian distribution such that
$\bm{y}_{1}^{T},\bm{y}_{2}^{T},\cdot,\bm{y}_{k}^{T}]^{T}\sim\mathcal{N}([f^{T}(\bm{\theta}_{1}),f^{T}(\bm{\theta}_{2}),\cdot,f^{T}(\bm{\theta}_{k})^{T}]^{T},\bm{\Omega}\otimes\bm{K})$ (25)
where $\bm{\Omega}\in\mathbb{R}^{k\times k}$ with $\bm{\Omega}_{i,j}=\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$. Then, a pGP on $\mathcal{G}(r,n)$ can be defined.
Projected Gaussian Process. A pGP on $\mathcal{G}(r,n)$ is defined by
$g\sim\mathcal{PGP}(m,\omega,\bm{K})$ (26)
@@
$\begin{pmatrix}p_{*}\\ \bm{p}^{\text{train}}\end{pmatrix}\sim\mathcal{N}_{\mathcal{G}_{*}\times\mathcal{G}_{1}\times\cdots\times\mathcal{G}_{k}}\left(\begin{pmatrix}m_{*}\\ \bm{m}^{\text{train}}\end{pmatrix},\begin{pmatrix}\bm{K}_{**},\bm{K}_{*}\\ \bm{K}_{*}^{T},\bm{K}^{\text{train}}\end{pmatrix}\right)$ (27)
@@
$)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}$)) (28)
@@
$\bm{u}_{*}=\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}((\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T},(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T},\cdots,(\mathcal{P}_{m(\bm{\theta}_{d})}^{-1}(p_{k}))^{T})^{T}$ (29)
@@
$\tilde{\bm{K}}_{*}=\bm{K}_{**}-\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}\bm{K}_{*}^{T}.$ (30)
[…]
@@
Figure 4 illustrates the process of predicting the optimal subspace using pGP. In this illustration, suppose that we obtain the snapshot data from three parameter settings $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$, as well as the optimal subspaces $p_{1}$, $p_{2}$ and $p_{3}$ spanned by the POD modes found at $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$. The prediction of the optimal subspace involves the following steps:
@@
[…]
(iv) Finally, the Gaussian distribution of $\bm{y}_{*}$ on $\mathbb{R}^{nr-r}$ (or, the predicted column vector $\bm{u}_{*}$ if only a point prediction is needed) is projected back to the Grassmann manifold through the projector $\mathcal{P}_{m(\bm{\theta}_{*})}$.
@@ 3.4 Discussions on model specification and parameter estimation
A close examination of (28) reveals an important difference between how predictions are made by conventional GP and the proposed pGP. Using a conventional GP, the point predictor is a linear combination of the observations and the weights are determined by the covariance function (i.e., the point prediction does not depend on the mean of the GP; a property well-known in the literature). For the proposed pGP, as indicated by (28), the point prediction does depend on the function $m(\bm{\theta})$ that describes the relationship between the parameter and basepoint on the Grassmann manifold. In other words, the choice of $m(\bm{\theta})$ determines the mapping $\mathcal{P}_{m}$. Two approaches can be used to determine the function $m(\bm{\theta})$. The first approach is to fix the basepoint for all parameters, i.e., $m(\bm{\theta})=\mu$. which can be obtained using a global POD basis. The second approach is to use another regression model to determine $m(\bm{\theta})$, such as the tree-based method described in Liu and Liu, (2023).
@@
In addition, because $n$ is usually large, a practical way to parameterize the $(nr-r)\times(nr-r)$ covariance matrix $\bm{K}$ is to let $\bm{K}=\sigma_{\bm{K}}\bm{I}$ and estimate $\sigma_{\bm{K}}$ from data. As for the kernel function $\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$, existing Gaussian, Exponential or other kernel functions can be used, and the parameter $\bm{\xi}$ can be estimated from data. Hence, once $m(\bm{\theta})$ is determined, $(\sigma_{\bm{K}},\bm{\xi})$ can be estimated using existing methods such as the cross-validation-based approach or maximum likelihood estimation described in Rasmussen and Williams, (2005).
@@
Also note that, for systems with high-dimensional parameter spaces (i.e., relatively large $d$), it is often the case that not all parameters are equally important in terms of their predictive power of the POD modes. In other words, the POD modes can be sensitive to some parameters while independent of the others. Hence, when constructing the statistical learning model that establishes the mapping from parameter space to Grassmann manifold, it is useful to identify those relevant parameters. To illustrate the idea, consider the squared Exponential kernel given by
$\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}$ (31)
@@
where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2},\cdots,\xi_{d+2}^{-2}\}$. Here, the parameters $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ play the role of characteristic length-scales for parameter $\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{d}$. If the estimated length-scales are very large, the corresponding parameters have little effects on the prediction and are practically removed from the model. This is known as the Automatic Relevance Determination (ARD) (Rasmussen and Williams,, 2005) and more details can be found in Sec.4.
@@
Finally, it is worth noting that the proposed pGP naturally involves the use of the global POD basis as its special case. Take the squared Exponential kernel above as an example. If $\xi_{1}=0$ and the estimated $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ are all close to zero, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ becomes practically zero. In this case, (29) immediately implies that $\bm{u}_{*}=\bm{0}$ and $)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}$)). If the global POD basis is chosen as the reference point (which spans a subspace in the Grassmann manifold where the tangent space is attached to) and the basepoint function $\bm{m}$ is invariant with $\bm{\theta}$, then, any subspace $p_{i}$, corresponding to the parameter $\bm{\theta}_{i}$, on the Grassmann manifold can be seen as i.i.d. samples from a pGD, $)\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}$)). Hence, the point prediction, $\hat{p}_{*}$, is just the subspace spanned by the global POD basis. More details can be found in Sec.4.